OpenStudy (anonymous):

-2 race to the power -1 ..??

OpenStudy (alexandervonhumboldt2):

???????

OpenStudy (alexandervonhumboldt2):

what do you mean?

OpenStudy (anonymous):

-2^-1??

OpenStudy (alexandervonhumboldt2):

-2^-1=1/-2=-1/2

OpenStudy (anonymous):

-2^-1=1/-2=-1/2=-0.5